# Point-hyperplane incidences via extremal graph theory

### v1

@@ Abstract
The study of counting point-hyperplane incidences in the $d$-dimensional space was initiated in the 1990’s by Chazelle and became one of the central problems in discrete geometry. It has interesting connections to many other topics, such as additive combinatorics and theoretical computer science. Assuming a standard nondegeneracy condition, i.e., that no $s$ points are contained in the intersection of $s$ hyperplanes, the currently best known upper bound on the number of incidences of $n$ points and $m$ hyperplanes in $\mathbb{R}^{d}$ is
@@
This bound by Apfelbaum and Sharir is based on geometrical space partitioning techniques.
In this paper, we propose a novel combinatorial approach to estimate the number of point-hyperplane incidences over arbitrary fields using forbidden induced patterns in incidence graphs. Perhaps surprisingly, this approach matches the best known bounds in $\mathbb{R}^{d}$ for many interesting values of $m,n,d$, e.g. when $m=n$ and $d$ is odd. Moreover, in finite fields our bounds are sharp as a function of $m$ and $n$ in every dimension. We also study the size of the largest complete bipartite graph in point-hyperplane incidence graphs with a given number of edges and obtain optimal bounds as well.
@@ 1 Introduction
The Szemerédi-Trotter theorem [30] is a fundamental result in combinatorial geometry, giving a sharp upper bound on the number of incidences in point-line configurations. It states that $n$ points and $m$ lines on the real plane determine at most $O((mn)^{2/3}+m+n)$ incidences, and as Erdős showed (see e.g. [11]), this bound is the best possible. This deep result found numerous applications and inspired a large number of generalizations and extensions over the last decades. For example, the problem of counting point-line incidences in $\mathbb{R}^{3}$ under certain non-degeneracy conditions has been studied by Guth and Katz [17] in their solution to the Erdős distinct distances problem. The Szemerédi-Trotter theorem has also been applied by Elekes [12] to derive sum-product estimates over the reals. For these and other applications of incidence theorems, we refer the reader to the survey [9].
@@
Extending incidence results to finite fields is more tricky, since standard space partitioning techniques do not apply anymore. The importance of incidence bounds in $\mathbb{F}_{p}^{2}$ stems from their close connection to sum-product estimates, as shown in the pioneering work of Bourgain, Katz and Tao [4]. Following this work, point-line incidence bounds over finite fields have been extensively studied, see e.g. [18, 22, 27].
@@
In this paper, we consider one of the most natural incidence questions in higher dimensions, which is to find the maximum number of incidences between $n$ points and $m$ hyperplanes in $\mathbb{F}^{d}$, for an arbitrary field $\mathbb{F}$. In dimension 3 and above, it is possible that all $m$ hyperplanes contain all $n$ points: take $n$ points on a line, and $m$ hyperplanes containing this line. In order to avoid such trivialities, it is standard to impose the further condition that the incidence graph is $K_{s,s}$-free, where $K_{s,s}$ denotes the complete bipartite graph with vertex classes of size $s$. Here, we think of $s$ as a large constant, which may depend on $d$, but no other parameters. Given a set of points $\mathcal{P}$ and set of hyperplanes $\mathcal{H}$, the incidence graph of $(\mathcal{P},\mathcal{H})$ is the bipartite graph $G(\mathcal{P},\mathcal{H})$ with vertex classes $\mathcal{P}$ and $\mathcal{H}$, and $\{x,H\}\subset\mathcal{P}\cup\mathcal{H}$ is an edge if $x\in H$. We denote by $I(\mathcal{P},\mathcal{H})$ the number of edges of the incidence graph, i.e. $I(\mathcal{P},\mathcal{H})$ is the number of incidences between $\mathcal{P}$ and $\mathcal{H}$.
1.1 Incidence bounds in higher dimensions
@@
The problem of bounding the number of edges in $K_{s,s}$-free incidence graphs has a long history, with the initial motivation coming from computational geometry. How compactly can one record the incidences in a configuration of points and hyperplanes? Instead of writing down whether each of the $mn$ pairs $(x,H)\in\mathcal{P}\times\mathcal{H}$ forms an incidence, one can compress the incidence graph by representing it as the union of few complete balanced bipartite subgraphs. Although this cannot be done very efficiently in general graphs (i.e., one needs $\Omega(\frac{n^{2}}{\log n})$ complete balanced bipartite subgraphs to represent a random balanced bipartite graph on $2n$ vertices with high probability), for incidence graphs Chazelle [6] developed a space partitioning technique which can be used to show that $O_{d,s}\big{(}n^{2-\frac{2}{d+1}}\log n\big{)}$ complete bipartite graphs suffice. Another motivation comes from the counting version of Hopcroft’s problem, which asks, given $n$ points and $m$ hyperplanes in $\mathbb{R}^{d}$, how fast can one determine the number of incidences. The work of Chazelle [6] achieved the first subquadratic algorithm for this problem and Erickson [13] gave a lower bound for the running time of algorithms classified as partitioning algorithms. He showed that particularly hard instances for determining the number of incidences are configurations $(\mathcal{P},\mathcal{H})$ without a compact representation, i.e. a configuration without a large complete bipartite graph in the incidence graph.
@@
Improving the works of Chazelle [6] and Brass and Knauer [5], Apfelbaum and Sharir [2] proved that if $\mathcal{P}$ is a set of $n$ points and $\mathcal{H}$ is a set of $m$ hyperplanes in $\mathbb{R}^{d}$, whose incidence graph contains no $K_{s,s}$, then
@@
This upper bound has not been improved in the past twenty years, however, matching lower bounds are only known for $d=2$, which coincides with the Szemerédi-Trotter theorem. The currently best known lower bound for $d\geq 3$ was recently achieved by Sudakov and Tomon [29], improving on the constructions of Brass and Knauer [5] and Balko, Cibulka, and Valtr [3]: if $s$ is sufficiently large with respect to $d$, then there exist a set of $n$ points $\mathcal{P}$ and a set of $m$ hyperplanes $\mathcal{H}$ in $\mathbb{R}^{d}$ whose incidence graph is $K_{s,s}$-free and
@@
Generalizations of the upper bound in (1) are also established (with an extra $o(1)$ term in the exponent) for incidences of points and semi-algebraic sets in $\mathbb{R}^{d}$, see [14]. The common feature in all of the proofs of (1) and related results is that they rely on certain space partitioning results such as cuttings [6, 7] or polynomial partitioning [17, 26]. These techniques are highly geometric and have no analogues over finite fields.
Another natural question asks about the maximum number of incidences between points and hyperplanes in a $d$-dimensional space $\mathbb{F}^{d}$ for any field $\mathbb{F}$. If $d=2$ and $p$ is a prime, taking every point of $\mathbb{F}_{p}^{2}$ and $m<p^{2}$ arbitrary lines, we get $n=p^{2}$ points, and $mp=\Theta(mn^{1/2})$ incidences, beating the Szemerédi-Trotter bound as long as $m\gg p$. On the other hand, a simple application of the Kővári-Sós-Turán theorem [19] shows that one cannot have more than $O(mn^{1/2}+n)$ incidences over any field, by noting that the incidence graph of points and lines is always $K_{2,2}$-free. As we mentioned already, the work of Bourgain, Katz, and Tao [4] shows that better upper bounds can be obtained over $\mathbb{F}_{p}$ assuming $m,n\leq p^{2-\delta}$.
However, point-hyperplane incidence bounds in higher dimensions turn out to be more elusive. One such result in three dimensions was obtained by Rudnev [24] (see also de Zeeuw [31] for a shorter proof). He proved that in $\mathbb{F}_{p}^{3}$, if no $s$ points are on a line, then there are at most $O(m\sqrt{n}+ms)$ incidences, assuming $n\leq p^{2}$. This point-plane incidence bound in $\mathbb{F}_{p}^{3}$ was later used to show various improved sum-product estimates over finite fields in [1, 22, 23]. For $d>3$, barely anything is known about incidences of points and hyperplanes in $\mathbb{F}_{p}^{d}$.
1.2 Sparse incidence graphs in vector spaces
In this paper, we propose a novel approach to derive bounds on the maximal number of incidences between $n$ points and $m$ hyperplanes in a $d$-dimensional vector space $\mathbb{F}^{d}$, assuming the incidence graph is $K_{s,s}$-free. We use almost purely graph-theoretic techniques and prove bounds which are sharp for the whole range of parameters $m$ and $n$ with a suitable choice of field. Surprisingly, for many interesting pairs of values $(m,n)$, our upper bound matches the upper bound in (1). This is fairly unexpected, as all proofs of (1) rely on highly geometric techniques, while we employ only combinatorial ideas. To give a snippet of our most general result, we prove the following in the special case $m=n$.
@@ 1.2 Sparse incidence graphs in vector spaces
Let $d,s,m,n$ be positive integers and $\alpha>0$ such that $m=n^{\alpha}$, and let $\mathbb{F}$ be a field. Let $\mathcal{P}$ be a set of $n$ points and $\mathcal{H}$ be a set of $m$ hyperplanes in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free. Then
@@
$I(\mathcal{P},\mathcal{H})\leq\begin{cases}O_{s,d}(n)&\mbox{ if }\alpha\in[0,\beta_{1}],\\ O_{s,d}(m^{1-\frac{1}{t}}n)&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ O_{s,d}(mn^{1-\frac{1}{d+2-t}})&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ O_{s,d}(m)&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Moreover, these bounds are tight, i.e. for every $d$, there exist $s=s(d)$ and $c=c(d)$ such that the following hold. For every $\alpha>0$ and sufficiently large integer $n$, there exist a field $\mathbb{F}=\mathbb{F}(d,n,\alpha)$, a set of $n$ points $\mathcal{P}$ and a set of $m=\lfloor n^{\alpha}\rfloor$ hyperplanes $\mathcal{H}$ in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free and
@@
$I(\mathcal{P},\mathcal{H})\geq\begin{cases}n&\mbox{ if }\alpha\in[0,\beta_{1}],\\ cm^{1-\frac{1}{t}}n&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ cmn^{1-\frac{1}{d+2-t}}&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ m&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Let us compare our result with (1) in the non-trivial regime $\alpha\in[1/d,d]$. Interestingly, in case $\alpha=\beta_{t}=\frac{t}{d+1-t}$ for some integer $t\in\{1,\dots,d\}$, Theorem 1.2 shows that the number of incidences is $O_{s,d}((mn)^{1-1/(d+1)})$, which exactly matches (1). The other extreme is when $\alpha=\alpha_{t}=\frac{t}{d+2-t}$ for some $t\in\{2,\dots,d\}$, in which case we get $O_{s,d}((mn)^{1-1/(d+2)})$. For other values of $\alpha$, our upper bound is $O_{s,d}((mn)^{1-\gamma})$ for some $\frac{1}{d+2}\leq\gamma\leq\frac{1}{d+1}$. However, as the second half of the theorem shows, these upper bounds cannot be improved unless further assumptions on the field $\mathbb{F}$ are made.
1.3 Large complete bipartite graphs in dense incidence graphs
@@
We also study another well-known problem of finding large complete bipartite graphs in dense incidence graphs. The motivation behind studying this problem is the intuitive understanding that a large number of incidences between points and hyperplanes is always explained by the existence of large complete bipartite subgraphs in the incidence graph. To make things more precise, given a set of points $\mathcal{P}$ and set of hyperplanes $\mathcal{H}$, we define $\operatorname{rs}(\mathcal{P},\mathcal{H})$ as the maximum number of edges in a complete bipartite subgraph of the incidence graph of $(\mathcal{P},\mathcal{H})$, i.e. the maximum of $r\cdot s$ over all $r$ and $s$ such that $K_{r,s}$ is a subgraph of the incidence graph. We consider configurations of $n$ points and $m$ hyperplanes with $\varepsilon mn$ incidences.
@@
In the real $d$-dimensional space, Apfelbaum and Sharir [2] proved that $\operatorname{rs}(\mathcal{P},\mathcal{H})=\Omega_{d}(\varepsilon^{d-1}mn)$ for $\varepsilon>\Omega(n^{-1/(d-1)})$. Also, if $\varepsilon>\Omega((mn)^{-\frac{1}{d-1}})$, then there exist a set of points $\mathcal{P}$ and set of hyperplanes $\mathcal{H}$ in $\mathbb{R}^{d}$ with $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq O_{d}(\varepsilon^{\frac{d+1}{2}}mn)$. Note that in case $d=3$, the lower and upper bounds match for $\varepsilon>\Omega(n^{-1/2})$. Do [8] improved the lower bound in $\mathbb{R}^{4}$ and $\mathbb{R}^{5}$ for a large range of $m,n,\varepsilon$ to match the upper bound up to logarithmic factors. One might be also interested how the minimum of $\operatorname{rs}(\mathcal{P},\mathcal{H})$ depends on the dimension $d$ as well, which is a question motivated by certain variants of the celebrated log-rank conjecture of Lovász and Saks [20]. To this end, it follows from Fox, Pach, and Suk [15] that $\operatorname{rs}(\mathcal{P},\mathcal{H})>\varepsilon^{d+1}2^{-O(d\log d)}mn$. For constant $\varepsilon$, this is improved by a recent result of Singer and Sudan [25], who obtained an exponential dependence on $d$ in the lower bound $\operatorname{rs}(\mathcal{P},\mathcal{H})=\Omega(\varepsilon^{2d}mn/d)$. We refer the interested reader to [25] about the relationship of this problem and the log-rank conjecture.
@@
Here, we extend and improve some of these results by considering arbitrary fields, and obtain tight bounds in all cases. Following Apfelbaum and Sharir [2], we define $\operatorname{rs}_{d}(n,m,I)=\min\operatorname{rs}(\mathcal{P},\mathcal{H})$, where the minimum is taken over all fields $\mathbb{F}$, all sets $\mathcal{P}\subset\mathbb{F}^{d}$ of at most $n$ points, all sets $\mathcal{H}$ of at most $m$ hyperplanes in $\mathbb{F}^{d}$, that satisfy $I(\mathcal{P},\mathcal{H})\geq I$. Note that requiring that $\mathcal{P}$ is a set of at most $n$ points instead of exactly $n$ points makes no crucial difference, since one can always add points to the set $\mathcal{P}$ which are incident to no hyperplanes (if $\mathbb{F}$ is a finite field, one might need to pass to an extension of $\mathbb{F}$). A similar remark holds for hyperplanes.
@@
Let $d$ and $m\geq n$ be positive integers, $\varepsilon\in(0,1)$ and $I=\varepsilon mn$. Then
$\operatorname{rs}_{d}(m,n,I),\operatorname{rs}_{d}(n,m,I)=\begin{cases}\Theta_{d}(\varepsilon^{d-1}mn)&\mbox{ if }\varepsilon>100\max\{m^{-\frac{1}{d}},n^{-\frac{1}{d-1}}\},\\ \Theta_{d}(\varepsilon m)&\mbox{ if }\varepsilon<\frac{1}{4}\max\{m^{-\frac{1}{d}},n^{-\frac{1}{d-1}}\}.\end{cases}$
@@
In the regime $\varepsilon\gg\max\{m^{-\frac{1}{d}},n^{-\frac{1}{d-1}}\}$, the implied constant of the lower bound is only exponential in $d$. Therefore, our theorem matches the dependence on $\varepsilon$ coming from the result of Apfelbaum and Sharir [2], as well as the exponential dependence on the dimension obtained by Sudan and Singer [25].

### v2

# Incidence bounds via extremal graph theory

@@ Abstract
The study of counting point-hyperplane incidences in the $d$-dimensional space was initiated in the 1990’s by Chazelle and became one of the central problems in discrete geometry. It has interesting connections to many other topics, such as additive combinatorics and theoretical computer science. Assuming a standard non-degeneracy condition, i.e., that no $s$ points are contained in the intersection of $s$ hyperplanes, the currently best known upper bound on the number of incidences of $m$ points and $n$ hyperplanes in $\mathbb{R}^{d}$ is
@@
This bound by Apfelbaum and Sharir is based on geometrical space partitioning techniques, which apply only over the real numbers.
In this paper, we propose a novel combinatorial approach to study such incidence problems over arbitrary fields. Perhaps surprisingly, this approach matches the best known bounds for point-hyperplane incidences in $\mathbb{R}^{d}$ for many interesting values of $m,n,d$, e.g. when $m=n$ and $d$ is odd. Moreover, in finite fields our bounds are sharp as a function of $m$ and $n$ in every dimension. We also study the size of the largest complete bipartite graph in point-hyperplane incidence graphs with a given number of edges and obtain optimal bounds as well.
Additionally, we study point-variety incidences and unit-distance problem in finite fields, and give tight bounds for both problems under a similar non-degeneracy assumption. We also resolve Zarankiewicz type problems for algebraic graphs. Our proofs use tools such as induced Turán problems, VC-dimension theory, evasive sets and Hilbert polynomials. Also, we extend the celebrated result of Rónyai, Babai and Ganapathy on the number of zero-patterns of polynomials to the context of varieties, which might be of independent interest.
@@ 1 Introduction
The Szemerédi-Trotter theorem [52] is a fundamental result in combinatorial geometry, giving a sharp upper bound on the number of incidences in point-line configurations. It states that $m$ points and $n$ lines on the real plane determine at most $O((mn)^{2/3}+m+n)$ incidences, and as Erdős showed (see e.g. [17]), this bound is the best possible. This deep result found numerous applications and inspired a large number of generalizations and extensions over the last decades. For example, the problem of counting point-line incidences in $\mathbb{R}^{3}$ under certain non-degeneracy conditions has been studied by Guth and Katz [27] in their solution to the Erdős distinct distances problem. The Szemerédi-Trotter theorem has also been applied by Elekes [18] to derive sum-product estimates over the reals. For these and other applications of incidence theorems, we refer the reader to the survey [13].
@@
Extending incidence results to finite fields is more tricky, since standard space partitioning techniques do not apply anymore. The importance of point-line incidence bounds in $\mathbb{F}_{p}^{2}$ stems from their close connection to sum-product estimates, as shown in the pioneering work of Bourgain, Katz and Tao [5]. Following this work, point-line incidence bounds over finite fields have been extensively studied, see e.g. [31, 37, 49].
@@
In this paper, we consider several natural incidence questions in higher dimensions. The first question we address is to find the maximum number of incidences between $m$ points and $n$ hyperplanes in $\mathbb{F}^{d}$, for an arbitrary field $\mathbb{F}$. In dimension 3 and above, it is possible that all $n$ hyperplanes contain all $m$ points: take $m$ points on a line, and $n$ hyperplanes containing this line. In order to avoid such trivialities, it is standard to impose the further condition that the incidence graph is $K_{s,s}$-free, where $K_{s,s}$ denotes the complete bipartite graph with vertex classes of size $s$. Here, we think of $s$ as a large constant, which may depend on $d$, but no other parameters. In geometric terms, this condition says that there do not exist $s$ hyperplanes whose intersection contains more than $s$ points.
Given a set of points $\mathcal{P}$ and set of geometric objects (e.g. hyperplanes) $\mathcal{H}$, the incidence graph of $(\mathcal{P},\mathcal{H})$ is the bipartite graph $G(\mathcal{P},\mathcal{H})$ with vertex classes $\mathcal{P}$ and $\mathcal{H}$, and $\{x,H\}\subset\mathcal{P}\cup\mathcal{H}$ is an edge if $x\in H$. We denote by $I(\mathcal{P},\mathcal{H})$ the number of edges of the incidence graph, i.e. $I(\mathcal{P},\mathcal{H})$ is the number of incidences between $\mathcal{P}$ and $\mathcal{H}$.
Beyond point-hyperplane incidences, we obtain sharp bounds for several other well-studied incidence problems. We find the maximum number of incidences between points and $d$-dimensional varieties of bounded degree over arbitrary fields assuming the incidence graph is $K_{s,s}$-free. Then, we determine the maximum number of unit distances between $n$ points in $\mathbb{F}^{d}$, assuming the unit distance graph is $K_{s,s}$-free. Finally, we conclude with Zarankiewicz’s problem for so called algebraic graphs.
1.1 Point-hyperplane incidences
@@
The problem of bounding the number of edges in $K_{s,s}$-free point-hyperplane incidence graphs has a long history, with the initial motivation coming from computational geometry. How compactly can one record the incidences in a configuration of points and hyperplanes? Instead of writing down whether each of the $mn$ pairs $(x,H)\in\mathcal{P}\times\mathcal{H}$ forms an incidence, one can compress the incidence graph by representing it as the union of complete bipartite subgraphs. This cannot be done very efficiently in general, for example if $G=G_{n,1/2}$ is the Erdős-Rényi random graph, one needs a list of $\Omega(\frac{n^{2}}{\log n})$ vertices to encode any partition. However, for incidence graphs Chazelle [9] developed a space partitioning technique to show that this can be improved to $O_{d,s}\big{(}n^{2-\frac{2}{d+1}}\log n\big{)}$ when $m=n$. Another motivation comes from the counting version of Hopcroft’s problem, which asks, given $m$ points and $n$ hyperplanes in $\mathbb{R}^{d}$, how fast can one determine the number of incidences. The work of Chazelle [9] achieved the first subquadratic algorithm for this problem and Erickson [21] gave a lower bound for the running time of algorithms classified as partitioning algorithms. He showed that particularly hard instances for determining the number of incidences are configurations $(\mathcal{P},\mathcal{H})$ without a compact representation, i.e. a configuration without a large complete bipartite graph in the incidence graph.
@@
Improving the works of Chazelle [9] and Brass and Knauer [6], Apfelbaum and Sharir [2] proved that if $\mathcal{P}$ is a set of $m$ points and $\mathcal{H}$ is a set of $n$ hyperplanes in $\mathbb{R}^{d}$, whose incidence graph contains no $K_{s,s}$, then
@@
This upper bound has not been improved in the past twenty years, however, matching lower bounds are only known for $d=2$, which coincides with the Szemerédi-Trotter theorem. The common feature in all of the proofs of (1) and related results is that they rely on certain space partitioning results such as cuttings [9, 10] or polynomial partitioning [27, 47]. These techniques are highly geometric and have no analogues over finite fields.
The currently best known lower bound for $d\geq 3$ was recently achieved by Sudakov and Tomon [51], improving on the constructions of Brass and Knauer [6] and Balko, Cibulka, and Valtr [3]: if $s$ is sufficiently large with respect to $d$, then there exist a set of $m$ points $\mathcal{P}$ and a set of $n$ hyperplanes $\mathcal{H}$ in $\mathbb{R}^{d}$ whose incidence graph is $K_{s,s}$-free and
@@
A natural variant of this problem asks about the maximum number of incidences between points and hyperplanes in a $d$-dimensional space $\mathbb{F}^{d}$ for a general field $\mathbb{F}$. If $d=2$ and $p$ is a prime, taking every point of $\mathbb{F}_{p}^{2}$ and $n<p^{2}$ arbitrary lines, we get $m=p^{2}$ points, and $np=\Theta(nm^{1/2})$ incidences, beating the Szemerédi-Trotter bound as long as $n\gg p$. On the other hand, a simple application of the Kővári-Sós-Turán theorem [32] shows that one cannot have more than $O(nm^{1/2}+m)$ incidences over any field, by noting that the incidence graph of points and lines is always $K_{2,2}$-free. The work of Bourgain, Katz, and Tao [5] shows that better upper bounds can be obtained over $\mathbb{F}_{p}$ assuming $m,n\leq p^{2-\delta}$.
However, point-hyperplane incidence bounds in higher dimensions turn out to be more elusive. One such result in three dimensions was obtained by Rudnev [43] (see also de Zeeuw [57] for a shorter proof). He proved that in $\mathbb{F}_{p}^{3}$, if no $s$ planes contain the same line, then there are at most $O(m\sqrt{n}+ms)$ incidences, assuming $n\leq p^{2}$. This point-plane incidence bound in $\mathbb{F}_{p}^{3}$ was later used to show various improved sum-product estimates over finite fields in [1, 37, 41]. For $d>3$, barely anything is known about incidences of points and hyperplanes in $\mathbb{F}_{p}^{d}$.
In this paper, we propose a novel approach to derive bounds on the maximal number of incidences between $m$ points and $n$ hyperplanes in a $d$-dimensional vector space $\mathbb{F}^{d}$, assuming the incidence graph is $K_{s,s}$-free. We combine new graph-theoretic techniques with probabilistic arguments and prove bounds which are sharp for the whole range of parameters $m$ and $n$ with a suitable choice of field. Surprisingly, for many interesting pairs of values $(m,n)$, our upper bound matches the upper bound (1). This is fairly unexpected, as all proofs of (1) rely on highly geometric techniques, while we employ only combinatorial ideas. To give a snippet of our most general result, we prove the following in the special case $m=n$.
@@
Let $d,s,m,n$ be positive integers and $\alpha>0$ such that $n=m^{\alpha}$, and let $\mathbb{F}$ be a field. Let $\mathcal{P}$ be a set of $m$ points and $\mathcal{H}$ be a set of $n$ hyperplanes in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free. Then
@@
$I(\mathcal{P},\mathcal{H})\leq\begin{cases}O_{s,d}(m)&\mbox{ if }\alpha\in(0,\beta_{1}],\\ O_{s,d}(m^{1-\frac{1}{d+2-t}}n)&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ O_{s,d}(mn^{1-\frac{1}{t}})&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ O_{s,d}(n)&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Moreover, these bounds are tight, i.e. for every $d$, there exist $s=s(d)$ and $c=c(d)$ such that the following hold. For every $\alpha>0$ and sufficiently large integer $m$, there exist a field $\mathbb{F}=\mathbb{F}(d,m,\alpha)$, a set of $m$ points $\mathcal{P}$ and a set of $n=\lfloor m^{\alpha}\rfloor$ hyperplanes $\mathcal{H}$ in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free and
@@
$I(\mathcal{P},\mathcal{H})\geq\begin{cases}m&\mbox{ if }\alpha\in(0,\beta_{1}],\\ cm^{1-\frac{1}{d+2-t}}n&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ cmn^{1-\frac{1}{t}}&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ n&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Let us compare our result with (1) in the non-trivial regime $\alpha\in[1/d,d]$. Interestingly, in case $\alpha=\beta_{t}=\frac{t}{d+1-t}$ for some integer $t\in\{1,\dots,d\}$, Theorem 1.2 shows that the number of incidences is $O_{s,d}((mn)^{1-1/(d+1)})$, which exactly matches (1). The other extreme is when $\alpha=\alpha_{t}=\frac{t}{d+2-t}$ for some $t\in\{2,\dots,d\}$, in which case we get $O_{s,d}((mn)^{1-1/(d+2)})$. For other values of $\alpha$, our upper bound is $O_{s,d}((mn)^{1-\gamma})$ for some $\frac{1}{d+2}\leq\gamma\leq\frac{1}{d+1}$.
The second half of the theorem shows that these upper bounds cannot be improved unless further assumptions on the field $\mathbb{F}$ are made. To show this, we consider optimal constructions of so called subspace evasive sets [16, 51]. We follow the ideas of [3, 6, 51] to transform large subspace evasive sets into point-hyperplane configuration with many incidences and no large bipartite subgraphs.
@@
We also study another well-known problem about finding large complete bipartite graphs in dense incidence graphs. The motivation behind studying this problem is the intuitive understanding that a large number of incidences between points and hyperplanes is always explained by the existence of large complete bipartite subgraphs in the incidence graph. To make things more precise, given a set of points $\mathcal{P}$ and set of hyperplanes $\mathcal{H}$, we define $\operatorname{rs}(\mathcal{P},\mathcal{H})$ as the maximum number of edges in a complete bipartite subgraph of the incidence graph of $(\mathcal{P},\mathcal{H})$, i.e. the maximum of $r\cdot s$ over all $r$ and $s$ such that $K_{r,s}$ is a subgraph of the incidence graph. We consider configurations of $m$ points and $n$ hyperplanes with $\varepsilon mn$ incidences.
@@
In the real $d$-dimensional space, Apfelbaum and Sharir [2] proved that $\operatorname{rs}(\mathcal{P},\mathcal{H})=\Omega_{d}(\varepsilon^{d-1}mn)$ for $\varepsilon>\Omega(n^{-1/(d-1)})$. Also, if $\varepsilon>\Omega((mn)^{-\frac{1}{d-1}})$, then there exist a set of points $\mathcal{P}$ and set of hyperplanes $\mathcal{H}$ in $\mathbb{R}^{d}$ with $\operatorname{rs}(\mathcal{P},\mathcal{H})\leq O_{d}(\varepsilon^{\frac{d+1}{2}}mn)$. Note that in case $d=3$, the lower and upper bounds match for $\varepsilon>\Omega(n^{-1/2})$. Do [12] improved the lower bound in $\mathbb{R}^{4}$ and $\mathbb{R}^{5}$ for a large range of $m,n,\varepsilon$ to match the upper bound up to logarithmic factors. One might be also interested how the minimum of $\operatorname{rs}(\mathcal{P},\mathcal{H})$ depends on the dimension $d$ as well, which is a question motivated by certain variants of the celebrated log-rank conjecture of Lovász and Saks [35]. In this regime, it follows from Fox, Pach, and Suk [23] that $\operatorname{rs}(\mathcal{P},\mathcal{H})>\varepsilon^{d+1}2^{-O(d\log d)}mn$. For constant $\varepsilon$, this is improved by a recent result of Singer and Sudan [46], who obtained an exponential dependence on $d$ in the lower bound $\operatorname{rs}(\mathcal{P},\mathcal{H})=\Omega(\varepsilon^{2d}mn/d)$. We refer the interested reader to [46] about the relationship of this problem and the log-rank conjecture.
@@
Here, we extend and improve some of these results by considering arbitrary fields, and obtain tight bounds in all cases. Following Apfelbaum and Sharir [2], we define $\operatorname{rs}_{d}(n,m,I)=\min\operatorname{rs}(\mathcal{P},\mathcal{H})$, where the minimum is taken over all fields $\mathbb{F}$, all sets $\mathcal{P}\subset\mathbb{F}^{d}$ of at most $m$ points, all sets $\mathcal{H}$ of at most $n$ hyperplanes in $\mathbb{F}^{d}$, that satisfy $I(\mathcal{P},\mathcal{H})\geq I$. Note that requiring that $\mathcal{P}$ is a set of at most $m$ points instead of exactly $m$ points makes no crucial difference, since one can always add points to the set $\mathcal{P}$ which are incident to no hyperplanes (if $\mathbb{F}$ is a finite field, one might need to pass to an extension of $\mathbb{F}$). A similar remark holds for hyperplanes.
@@
Let $d$ and $m\leq n$ be positive integers, $\varepsilon\in(0,1)$ and $I=\varepsilon mn$. Then
$\operatorname{rs}_{d}(m,n,I),\operatorname{rs}_{d}(n,m,I)=\begin{cases}\Theta_{d}(\varepsilon^{d-1}mn)&\mbox{ if }\varepsilon>100\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\},\\ \Theta_{d}(\varepsilon n)&\mbox{ if }\varepsilon<\frac{1}{4}\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}.\end{cases}$
@@
In the regime $\varepsilon\gg\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}$, the implied constant of the lower bound is only exponential in $d$. Therefore, our theorem matches the dependence on $\varepsilon$ coming from the result of Apfelbaum and Sharir [2], as well as the exponential dependence on the dimension obtained by Sudan and Singer [46].